A 2.5 kg object is whirled in a vertical circle whose radius is 0.89 m. If the time of one revolution is 0.94 s, the magnitude of the tension in the string (assuming uniform speed) when it is at the bottom is ___x102N. (Give your answer with 2 sig digs, and do not include units)


1
Expert's answer
2021-07-14T10:15:13-0400
"T=F_c+F_g,""T=\dfrac{mv^2}{r}+mg,""T=\dfrac{m(\dfrac{2\pi r}{T})^2}{r}+mg,""T=\dfrac{4\pi^2rm}{T^2}+mg,""T=\dfrac{4\pi^2\cdot0.89\ m\cdot2.5\ kg}{(0.94\ s)^2}+2.5\ kg\cdot9.8\ \dfrac{m}{s^2}=1.2\cdot10^2\ N."